# A Survey of Sustainability in Large Language Models: Applications, Economics, and Challenges

### v1

@@ I-A Research Questions and Inquiries
The research questions explored in this paper, along with their corresponding inquiries, are as follows:
@@ III-A Environmental Concerns
  • Carbon Emissions: The energy-intensive processes involved in Large Language Models (LLMs) contribute significantly to carbon dioxide equivalent (CO2e) emissions. The article ”Double Jeopardy and Climate Impact in the Use of Large Language Models” highlights the carbon footprint arising from the training and operation of LLMs, emphasizing the importance of adopting sustainable AI practices [39]. Faiz et al. [37] introduced a framework, LLMCarbon (Fig. 1), to estimate emissions across the lifecycle of LLMs, including both operational and embodied carbon. Similarly, Luccioni et al. [40] analyzed the BLOOM model’s emissions (Table III), highlighting the statistics for equipement manufacturing, dynamic power consumption and idle energy consumption. To mitigate these emissions, strategies such as using energy-efficient hardware, implementing optimization techniques like pruning and quantization [41], and transitioning to renewable energy sources [42] are recommended. These approaches are critical for reducing LLMs’ environmental impact while enabling sustainable AI development.

### v2

@@ I-A Research Questions and Inquiries
The research questions explored in this paper (Figure 1), along with their corresponding inquiries, are as follows:
@@
{forest}
for tree=forked edges, grow’=0, draw, rounded corners, node options=align=center,, text width=2.5cm, level distance=1.5cm, sibling distance=1.8cm, , [Research Questions [Tasks and Domains of LLMs, fill=blue!20 [Identify Practical Applications, fill=blue!10] ] [Environmental and Economic Costs, fill=green!20 [Evaluate Training
and Deployment Costs, fill=green!10] [Assess Sustainability, fill=green!10] ] [Sustainable and Energy-Efficient LLMs, fill=yellow!20 [Explore Energy
Reduction Strategies, fill=yellow!10] ] ]
Figure 1: Hierarchical Tree of Research Questions and Objectives for LLM Sustainability
@@ II LLM Applications
Text-to-Text Generates text from text descriptions Education, Marketing, e-commerce.
@@
  • Text-to-Text: Text-to-Text Large Language Models (LLMs) enable the generation of text from textual descriptions, with applications such as text summarization[2, 3, 4, 5], multilingual text generation[6, 7, 8, 9], and question answering [10, 11, 12] using RAG[13]. These models are widely used for condensing information, enabling cross-lingual communication, and developing intelligent systems like chatbots and virtual assistants. Their ability to process and generate coherent, contextually appropriate text across domains demonstrates their transformative potential in automating complex language tasks.
@@ III-A Environmental Concerns
  • Carbon Emissions: The energy-intensive processes involved in Large Language Models (LLMs) contribute significantly to carbon dioxide equivalent (CO2e) emissions. Carbon emissions vary significantly by model size and training location, as shown in Figures 2 and 3. The article ”Double Jeopardy and Climate Impact in the Use of Large Language Models” highlights the carbon footprint arising from the training and operation of LLMs, emphasizing the importance of adopting sustainable AI practices [51]. Faiz et al. [49] introduced a framework, LLMCarbon (Fig. 4), to estimate emissions across the lifecycle of LLMs, including both operational and embodied carbon. Similarly, Luccioni et al. [52] analyzed the BLOOM model’s emissions (Table III), highlighting the statistics for equipement manufacturing, dynamic power consumption and idle energy consumption. To mitigate these emissions, strategies such as using energy-efficient hardware, implementing optimization techniques like pruning and quantization [53], and transitioning to renewable energy sources [54] are recommended. These approaches are critical for reducing LLMs’ environmental impact while enabling sustainable AI development.
@@
Figure 2: Carbon emissions for training BERT (8x V100s, 36 hrs): highest in Central US/Australia, lowest in Norway/France. From Dodge et al. [59]
Figure 3: Carbon emissions for training experiments: larger models emit significantly more than smaller ones. From Dodge et al.[59]